40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Feb 10 22:21:55 2014, manasa, Update, Green Locking, X and Y arm green tuned 
    Reply  Tue Feb 11 21:22:02 2014, manasa, Update, Green Locking, ALS X and Y arm restored 
       Reply  Tue Feb 11 22:17:06 2014, Koji, Update, Green Locking, ALS X and Y arm restored 
    Reply  Thu Feb 13 16:48:33 2014, Steve, Update, Green Locking, X and Y arm green tuned details 
Message ID: 9633     Entry time: Thu Feb 13 16:48:33 2014     In reply to: 9621
Author: Steve 
Type: Update 
Category: Green Locking 
Subject: X and Y arm green tuned details 

Quote:

Y arm green: Nothing much was disturbed. I touched the steering mirrors and brought GTRY from 0.2 to 0.9.

X arm green: The PDH lock was not very stable mostly because of the low power in green. I changed the oven temperature for the doubler to 36.4 corresponding to maximum green power. GTRX increased from 0.1 to 0.9

Both the X and Y arm green alignment were tuned on the PSL table to their respective beat PDs.

The PSL green shutter was not responding to the medm buttons. I found the PSL green shutter set to 'local' and 'N.O' (these are switches in the shutter controller). I do not see any elog and not sure as to why the controller was even touched in the first place. I set the shutter controls to 'remote' and 'N.C'.

 ETMX green power at shutter 3.6 mW at 36.35 C doubler crystal temp.   [ Innolight IR settings 2.0 A,  40.83 C,  500 mW before Faraday 1/2 plate ] 

 ETMY green power at shutter 0.75 mW  at 35.8 C  doubler crytal temp.  [ NPRO IR settings 1.82A,  231 mW_ display,  DT 21 C, DTEC +1V, LT 40 C, LTEC 0.1V, T +41.041 ]  

ELOG V3.1.3-